OpenStudy (anonymous):

complete following identity: 1/1+sinx + 1/1-sinx I'm so lost on this question.

OpenStudy (anonymous):

try adding \[\frac{1}{1+b}+\frac{1}{1-b}\] first to see what you get. then replace b by sine

OpenStudy (anonymous):

i will write it if you like

OpenStudy (anonymous):

any additional info would be appreciated as this blows my mind.

OpenStudy (anonymous):

what else would you do here?

OpenStudy (anonymous):

ok lets do the algebra first. usually these work like this 1) some algebra 2) a trig identity 3) some more algebra (maybe) so rather than cluttering up your paper with sines lets just add

OpenStudy (anonymous):

\[\frac{1}{1+b}+\frac{1}{1-b}=\frac{1-b+1+b}{(1+b)(1-b)}\]

OpenStudy (anonymous):

simplify to get \[\frac{2}{1-b^2}\]notice this part had nothing at all to do with trig

OpenStudy (anonymous):

now replace b by sine and write \[\frac{2}{1-\sin^2(x)}\] and make the one trig step here, namely recall that \[\cos^2(x)+\sin^2(x)=1\] so \[1-\sin^2(x)=\cos^2(x)\]

OpenStudy (anonymous):

you get \[\frac{2}{\cos^2(x)}=2\sec^2(x)\]

OpenStudy (anonymous):

thank you so much. this makes more sense than the tutors do for class.